Calculate the mass of salt (NaCl) that is expected to be made based upon the mass of NaHCO3 used. This is called the 'theoretical mass of NaCl'.
1
Expert's answer
2020-12-15T07:49:14-0500

M(NaCl) = 23+35.5 = 58.5 g/mol

M(NaHCO3) = 23+1+12+16*3 = 84 g/mol

As NaCl and NaHCO3 have the same amount of Na, theoretical mass of NaCl will be:

m(NaCl, theor) = m(NaHCO3)/M(NaHCO3)*M(NaCl) = m(NaHCO3)/84*58.5 = m(NaHCO3)/1.436

Answer: m(NaHCO3)/1.436